OpenStudy (anonymous):

okay if i have 30 nickels and dimes worth $2.40. how many nickles and dimes do i have ??

OpenStudy (lgbasallote):

wait...do you mean 30 nickels and 30 dimes that equal to 2.40? Or 30 coins altogether?

OpenStudy (anonymous):

together!!

OpenStudy (anonymous):

???????

OpenStudy (lgbasallote):

your problem is a little difficult :( i'm sorry i wasn't able to answer it.

OpenStudy (anonymous):

its okay

OpenStudy (lgbasallote):

let x = no. of nickels y = no. of dimes x + y = 30 5(cents)x + 10(cents)y = 240 multiply the first equation by -5 -5x - 5y = -150 5x + 10y = 240 5y = 90 y = 19 There are 19 dimes and 11 nickels. I remembered how to do it :D It seems i made a mistake in my solution before that's why I can't solve it.